4. (a) If a competitive industry faces an increase in demand, what are the steps by which a competitive market ensures increased output? (b) Assume that a competitive firms Total Cost of producing output q is given by TC(Q)-3 + 3q + q2 . Assume that the market price of the firms product is e. What level of output will the firm produce? What is the firms profit? Will the firm be earning a positive, negative, or zero profit in the short run? What would you expect to happen in the long run?
Answer #1

4 a. As there is an increase in demand, price increases which leads to an increase in profit. As a result , producers will increase the supply. This will lead to a new equilibrium with a higher quantity and price.

b. Tc= 3+3q+q^2

Marginal cost = 2q + 3

Price =€9

Firms will produce that level of output where

Marginal cost = Marginal revenue= price

2q+3 = 9

2q = 6

q = 3

Thus , firms will produce 3 units of output.

Firms profit = revenue - cost

Revenue = P *q

= 9*3

= € 27

Cost = 3 + 3*3 + 3^3

= 3+ 9 + 27

= € 39

Profit = 27- 39

= -12

Therefore, firm is earning negative profit in the short run.

In the long run , if the firm continues to incur losses, then it will exit from the industry.